OpenStudy (anonymous):

The sum of three consecutive odd integers is 33. What is the first integer? 7 9 11 13

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (raden):

let the first number is x, then x + (x+2) + (x+4) = 33

OpenStudy (experimentx):

(2n-1)+(2n+1)+(2n+3) = 33 2n - 1 =?

OpenStudy (anonymous):

which 1 do i go by

OpenStudy (experimentx):

find the value of n ... then find the value of 2n-1 your three consecutive integers is 2n-1, 2n+1, and 2n+3

OpenStudy (anonymous):

7

OpenStudy (experimentx):

RadEn methods works too ... but when you are dealing with odd number it's customary to write 2n+1 or 2n-1.

OpenStudy (raden):

oh, sorry .. i was wrong make the model maths to solve the problem

OpenStudy (anonymous):

what is the answer yall are confusing me

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%282n-1%29%2B%282n%2B1%29%2B%282n%2B3%29+%3D+33 2n-1=10-1=9 so the numbers are 9,11, 13

OpenStudy (raden):

yea, that's right 9+11+13 = 33 hehe..